# Revisiting Stochastic Proximal Point Methods: Generalized Smoothness and Similarity

### v1

@@ 1 Introduction
(Bauschke et al., 2017). Proximal methods, which leverage proxs of functions (Parikh & Boyd, 2014; Condat et al., 2023b), are known for their robustness and resilience to the choice of stepsize, often allowing for the use of larger stepsizes compared to standard gradient-based methods. Ryu & Boyd (2014) provide convergence rate guarantees for SPPM and emphasize its stability with respect to inaccuracies in learning rate selection; a property not typically observed in SGD. Asi & Duchi (2019) investigate a more general framework, AProx, which encompasses SPPM as a special case. They establish both stability and convergence rate results for AProx under convexity assumptions. Moreover, SPPM has been shown to achieve convergence rates comparable to those of SGD across a variety of algorithmic settings (Richtárik et al., 2024).
@@
  • New generalized smoothness assumption, called $\boldsymbol{\phi}$-smoothness. We investigate the most general conditions required for the convergence of SPPM and introduce the novel notion of $\phi$-smoothness (Section 4). Under this assumption, we establish rigorous convergence guarantees and explore various special cases, highlighting the specific effects and implications of the proposed framework.
@@ 2 Preliminaries
where $[x,y]$ is the interval connecting the points $x$ and $y$.
@@
Then we consider the expected similarity assumption. Initially, the concept of similarity was introduced for twice-differentiable functions and is commonly referred to as Hessian similarity (Tian et al., 2022). For functions that are only once differentiable, analogous assumptions have been employed in the literature. In this work, we propose an even more general assumption introduced by Sadiev et al. (2024). We call it Star Similarity since it only has to hold with respect to a solution $x_{\star}$.
@@ 3 Related work
Stochastic Gradient Descent (SGD) (Robbins & Monro, 1951) is a fundamental and widely used optimization algorithm for training machine learning models. Due to its efficiency and scalability, it has become the backbone of modern deep learning, with state-of-the-art training methods relying on various adaptations of SGD (Zhou et al., 2020; Wu et al., 2020; Sun, 2020). Over the years, the algorithm has been extensively studied, leading to a deeper understanding of its convergence properties, robustness, and efficiency in different settings (Bottou et al., 2018; Gower et al., 2019; Khaled & Richtárik, 2020; Demidovich et al., 2023). This ongoing theoretical research continues to refine SGD and its variants, ensuring their effectiveness in large-scale and complex learning tasks (Carratino et al., 2018; Nguyen et al., 2018; Yang et al., 2024). Notably, methods designed to leverage the smoothness of the objective often struggle in Deep Learning, where optimization problems are inherently non-smooth. For instance, while variance-reduced methods (Johnson & Zhang, 2013; Defazio et al., 2014; Schmidt et al., 2017; Nguyen et al., 2017; Malinovsky et al., 2023; Cai et al., 2023) theoretically offer faster convergence for finite sums of smooth functions, they are often outperformed in practice by standard, non-variance-reduced methods (Defazio & Bottou, 2019). These challenges highlight the need to explore alternative assumptions that go beyond the standard smoothness assumption.
@@
Stochastic proximal point methods (Bertsekas, 2011) have been extensively studied across different settings due to their versatility and strong theoretical properties. This framework can encompass various optimization algorithms, making it a unifying approach for analyzing and designing new methods. One of its key advantages is enhanced stability, which helps mitigate the challenges of variance in stochastic optimization. Additionally, it is particularly well-suited for non-smooth problems, where traditional smoothness-based methods may struggle. (Bianchi, 2016; Toulis et al., 2016; Davis & Drusvyatskiy, 2019) These properties make stochastic proximal point methods a valuable tool in both theoretical analysis and practical applications (Davis & Yin, 2017; Condat et al., 2022; Condat & Richtárik, 2024).
Stochastic proximal methods have become increasingly important in Federated Learning due to their ability to handle decentralized optimization problems efficiently (Konečnỳ, 2016). Some researchers propose replacing the standard local update steps with the proximity operator, which provides a more robust framework for understanding the behavior of local methods and can lead to faster convergence rates by improving the optimization process (Li et al., 2020; T Dinh et al., 2020; Jhunjhunwala et al., 2023; Grudzień et al., 2023b; Li et al., 2024a). On the other hand, other studies focus on interpreting the aggregation step as a proximity operator, which allows for a more efficient combination of local updates (Mishchenko et al., 2022; Condat & Richtárik, 2022; Malinovsky et al., 2022; Condat et al., 2023a; Hu & Huang, 2023; Jiang et al., 2024a).
@@ 5 Convergence Results under $\phi$-Smoothness
where $\hat{x}^{k}$ is a vector chosen from the collection of iterates $x_{0}$, …, $x_{k-1}$ uniformly at random.
@@
It is worth noting that if the prox is computed exactly, the additional gradient step becomes redundant. In this case, $\nabla f_{\xi}(\hat{x})=\frac{1}{\gamma}(x_{k}-\hat{x}_{k})$, and we have $x_{k+1}=\mathrm{prox}_{\gamma f_{\xi}}(x_{k})$. SPPM with inexact prox (SPPM-inexact) is shown as Algorithm 2.
@@ 6 Convergence Results under Expected Similarity
6.1 Interpolation Regime
In this section we discuss the convergence results under expected similarity (Assumption 2.8) and interpolation regime (Assumption 2.4). The main concept of the proof is to introduce the average iterate
@@ 6.1 Interpolation Regime
where we denote by $\mathcal{F}_{k}$ the $\sigma$-algebra generated by the collection of random variables up to iteration $k$. This approach enables us to derive the complexity of the algorithm. Since we utilize the expected similarity assumption, we do not require any form of smoothness.
We analyze two regimes under this framework. The first is the interpolation regime, as defined in Assumption 2.4. In this setting, linear convergence to the exact solution is achieved, because convergence is not hindered by randomness.
Theorem 6.1.
Let Assumptions 2.1 (Differentiability), 2.2 (Convexity), 2.3 (Strong convexity of $f$), 2.4 (Interpolation), and 2.8 (Star Similarity) hold. If the stepsize satisfies $\gamma\leq\frac{\mu}{2\delta_{*}^{2}}$, then the iterates of SPPM satisfy, for every $k\geq 0$,
$\mathbb{E}\big{[}\|x_{k}-x_{*}\|^{2}\big{]}\leq\left(1-\min\left(\frac{\gamma\mu}{4},\frac{1}{2}\right)\right)^{k}\|x_{0}-x_{*}\|^{2}.$
It is worth noting that, in our result, the contraction factor is independent of the distance $\|x_{0}-x_{*}\|^{2}$ between the starting point and the solution, unlike in the case of $\phi$-smoothness. However, in this setting, the stepsize must satisfy the condition $\gamma\leq\frac{\mu}{2\delta_{*}^{2}}$.
Next, we establish convergence guarantees for SPPM-inexact. In this case, the contraction factor is not merely scaled by a constant but takes the form of the minimum of two terms. The first term, $\frac{\gamma\mu}{4}$, corresponds to the contraction factor in the exact version, differing only by a constant factor of 2. The second term, $\frac{1-c}{2}$, depends on the approximation constant $c$.
Theorem 6.2.
Let Assumptions 2.1 (Differentiability), 2.2 (Convexity), 2.4 (Interpolation) and 2.8 (Star Similarity) hold. Consider SPPM-inexact with Assumption 2.9 satsified. If the stepsize satisfies $\gamma\leq\frac{\mu}{4\delta_{*}^{2}}$ and $T$ is chosen sufficiently large such that $\frac{\eta\gamma^{2}}{T^{\alpha}}\leq c<1$, then the iterates of SPPM-inexact satisfy, for every $k\geq 0$,
$\mathbb{E}\left[\|x_{k}-x_{*}\|^{2}\right]\leq\left(1\!-\!\min\left(\frac{\gamma\mu}{4},\frac{1\!-\!c}{2}\right)\right)^{k}\|x_{0}-x_{*}\|^{2}.$
It is important to note that the stepsize bound depends on the strong convexity parameter $\mu$ and the similarity constant $\delta_{*}$, which are often difficult to estimate in practice due to their reliance on problem-specific properties. Despite this challenge, avoiding such a restriction seems unlikely, as these parameters are crucial to the stability and convergence of the method. Developing adaptive strategies to estimate or reduce dependence on these constants could be a valuable direction for future research.
6.2 Non-Interpolation Regime
In this section, we derive results without assuming that all stochastic gradients vanish at the solution. In this setting, convergence is impacted by the variance, defined as follows.
@@ 6.2 Non-Interpolation Regime
The assumption stated above is very mild and satisfied in virtually every realistic scenario. First, we present the convergence result for the exact formulation of SPPM.
@@
Let Assumptions 2.1 (Differentiability), 2.2 (Convexity), 2.3 (Strong convexity of $f$), 2.8 (Star Similarity) and 6.3 hold. If the stepsize satisfies $\gamma\leq\frac{\mu}{4\delta_{*}^{2}}$, then SPPM satisfies, for every $k\geq 0$,
@@
The result shows linear convergence to a neighborhood of the solution, with the neighborhood size proportional to the variance $\sigma_{*}^{2}$ from Assumption 6.3, and dependent on the stepsize $\gamma$. Using a decaying stepsize schedule leads to sublinear convergence to the exact solution.
@@
Let Assumptions 2.1 (Differentiability), 2.2 (Convexity), 2.8 (Star Similarity) and 6.3 (Bounded Variance) hold. Consider SPPM-inexact with Assumption 2.9 satisfied. If the stepsize satisfies $\gamma\leq\frac{\mu(1-c)}{4\delta_{*}^{2}}$ and $T$ is chosen sufficiently large such that $\frac{\eta\gamma^{2}}{T^{\alpha}}\leq c$, where $10<c<$ 1 is a constant, then SPPM-inexact satisfies, for every $k\geq 0$,
$\mathbb{E}\big{[}\|x_{k}-x_{*}\|^{2}\big{]}$ $\leq\left(1-\frac{1}{4}\min\left(\gamma\mu,1\right)\right)^{k}\|x_{0}-x_{*}\|^{2}+\max\left(\frac{1}{\gamma\mu},1\right)\frac{8\gamma^{2}\sigma_{*}^{2}}{(1-c)}.$
The result guarantees convergence to a neighborhood of the solution, with the neighborhood size increasing by a factor of $\frac{1}{1-c}$, where $c$ is the approximation constant. Although the contraction factor remains unchanged, the stepsize must be reduced by a factor of $(1-c)$ to ensure convergence. This reduction may slow convergence compared to the exact version of SPPM, but the method remains effective as long as the approximation error is controlled.
@@ 7 Experiments
In this section, we present numerical experiments conducted for the optimization problem in the finite-sum form (2), where the functions $f_{i}$ take the specific form $f_{i}(x)=a_{i}\|x\|^{2s},$ with an integer $s\geq 2$ and $a_{i}\in\mathbb{R}^{+}$ for all $i=1,\dots,n$. Each function $f_{i}$ is $(L_{0},L_{1})$-smooth, with $L_{0}=2s$ and $L_{1}=2s-1$, and is convex. We conduct three sets of experiments to analyze the behavior of SPPM-inexact under different conditions.
@@
The first set of experiments investigates the impact of different stepsizes on the convergence of the proposed method. Specifically, we aim to demonstrate that the method converges for any positive stepsize, provided that an appropriately chosen tolerance for the inner solver is used, and that larger stepsizes lead to faster convergence rates. To validate this, we analyze three different values of $s$, namely $s=2,3,4$. For each case, five different stepsizes are tested: $\gamma=0.1,1,10,100,1000$. The number of functions is set to $n=1000$, the dimension to $d=100$, and an inexact solver is employed with a stopping criterion based on the squared norm of the gradient, with an accuracy threshold of $10^{-12}$: $\|\nabla\Psi_{k}(\hat{x}_{k})\|^{2}\leq 10^{-12}.$ As shown in Figure 1, the method converges for all chosen stepsizes. While larger stepsizes accelerate convergence in terms of the number of iterations, they also increase the computational cost per iteration. These observations confirm our theoretical findings.

### v2

@@ 1 Introduction
(Bauschke & Combettes, 2017). Proximal methods, which leverage proxs of functions (Parikh & Boyd, 2014; Condat et al., 2023b), are known for their robustness and resilience to the choice of stepsize, often allowing for the use of larger stepsizes compared to standard gradient-based methods. Ryu & Boyd (2014) provide convergence rate guarantees for SPPM and emphasize its stability with respect to inaccuracies in learning rate selection; a property not typically observed in SGD. Asi & Duchi (2019) investigate a more general framework, AProx, which encompasses SPPM as a special case. They establish both stability and convergence rate results for AProx under convexity assumptions. Moreover, SPPM has been shown to achieve convergence rates comparable to those of SGD across a variety of algorithmic settings (Richtárik et al., 2024).
@@
  • New generalized smoothness assumption, called $\boldsymbol{\phi}$-smoothness. We investigate more general conditions under which SPPM converges and introduce the novel notion of $\phi$-smoothness (Section 4). Under this assumption, we establish rigorous convergence guarantees and explore various special cases, highlighting the specific effects and implications of the proposed framework.
@@ 2 Preliminaries
where $[x,y]$ denotes the line segment connecting points $x$ and $y$.
@@
Then we consider the expected similarity assumption. Initially, the concept of similarity was introduced for twice-differentiable functions and is commonly referred to as Hessian similarity (Tian et al., 2022). For functions that are only once differentiable, analogous assumptions have been employed in the literature. In this work, we consider an even more general assumption introduced by Sadiev et al. (2024). We call it Star Similarity since it only has to hold with respect to a solution $x_{\star}$.
@@ 3 Related work
Stochastic Gradient Descent (SGD) (Robbins & Monro, 1951) is a fundamental and widely used optimization algorithm for training machine learning models. Due to its efficiency and scalability, it has become the backbone of modern deep learning, with state-of-the-art training methods relying on various adaptations of SGD (Zhou et al., 2020; Wu et al., 2020; Sun, 2020). Over the years, the algorithm has been extensively studied, leading to a deeper understanding of its convergence properties, robustness, and efficiency in different settings (Bottou et al., 2018; Gower et al., 2019; Khaled & Richtárik, 2020; Demidovich et al., 2023). This ongoing theoretical research continues to refine SGD and its variants, ensuring their effectiveness in large-scale and complex learning tasks (Carratino et al., 2018; Nguyen et al., 2018; Yang et al., 2024). Notably, methods designed to leverage the smoothness of the objective often struggle in Deep Learning, where optimization problems are inherently nonsmooth. For instance, while variance-reduced methods (Johnson & Zhang, 2013; Defazio et al., 2014; Schmidt et al., 2017; Nguyen et al., 2017; Malinovsky et al., 2023; Cai et al., 2023) theoretically offer faster convergence for finite sums of smooth functions, they are often outperformed in practice by standard, non-variance-reduced methods (Defazio & Bottou, 2019). These challenges highlight the need to explore alternative assumptions that go beyond the standard smoothness assumption.
@@
Stochastic proximal point methods (Bertsekas, 2011) have been extensively studied across different settings due to their versatility and strong theoretical properties. This framework can encompass various optimization algorithms, making it a unifying approach for analyzing and designing new methods. One of its key advantages is enhanced stability, which helps mitigate the challenges of variance in stochastic optimization. Additionally, it is particularly well-suited for nonsmooth problems, where traditional smoothness-based methods may struggle. (Bianchi, 2016; Toulis et al., 2016; Davis & Drusvyatskiy, 2019) These properties make stochastic proximal point methods a valuable tool in both theoretical analysis and practical applications (Davis & Yin, 2017; Condat et al., 2022; Condat & Richtárik, 2024).
Stochastic proximal methods have become increasingly important in Federated Learning due to their ability to handle decentralized optimization problems efficiently (Konečnỳ, 2016). Some researchers propose replacing the standard local update steps with the proximity operator, which provides a more robust framework for understanding the behavior of local methods and can lead to faster convergence rates by improving the optimization process (Li et al., 2020; T Dinh et al., 2020; Jhunjhunwala et al., 2023; Grudzień et al., 2023b; Li et al., 2024a). On the other hand, other studies focus on interpreting the aggregation step as a proximity operator, which allows for a more efficient combination of local updates (Mishchenko et al., 2022; Condat & Richtárik, 2023; Malinovsky et al., 2022; Condat et al., 2023a; Hu & Huang, 2023; Jiang et al., 2024a).
@@ 5 Convergence Results under $\phi$-Smoothness
where $\hat{x}_{k}$ is a vector chosen from the collection of iterates $x_{0}$, …, $x_{k-1}$ uniformly at random.
@@
It is worth noting that if the prox is computed exactly, the additional gradient step becomes redundant. In this case, $\nabla f_{\xi}(\hat{x}_{k})=\frac{1}{\gamma}(x_{k}-\hat{x}_{k})$, and we have $x_{k+1}=\mathrm{prox}_{\gamma f_{\xi}}(x_{k})$. SPPM with inexact prox (SPPM-inexact) is shown as Algorithm 2.
@@
Remark 5.5 (Realizability of Assumption 2.9 by a parameter-free inner solver).
Although $\phi$-smoothness does not imply global Lipschitz continuity, the proximal subproblem $\Psi_{k}$ is $\tfrac{1}{\gamma}$–strongly convex and has a locally Lipschitz gradient on a bounded set containing the iterates (independent of $k$). Consequently, gradient descent on $\Psi_{k}$ with Armijo backtracking (requiring no knowledge of $L_{0}$, $L_{1}$, or explicit values of $\phi$) converges linearly. In particular, the inexactness model in Assumption 2.9 is realizable by such a parameter-free inner method. For more details, see Appendix B.2.1.
@@ 6 Convergence Results under Expected Similarity
In this section, we discuss the convergence results under expected similarity (Assumption 2.8) without assuming that all stochastic gradients vanish at the solution. The main concept of the proof is to introduce the average iterate:
@@
where we denote by $\mathcal{F}_{k}$ the $\sigma$-algebra generated by the randomness (e.g., stochastic gradients or iterates) up to iteration $k$. This approach enables us to derive the complexity of the algorithm. Since we utilize the expected similarity assumption, we do not require any form of smoothness.
In this setting, convergence is impacted by the variance, defined as follows:
@@
Remark 6.2.
The interpolation regime, where stochastic gradients vanish at the solution, is a special case of the general setting with $\sigma_{*}^{2}=0$. Detailed convergence results for this regime are provided in Appendix D.
First, we present the convergence result for the exact formulation of SPPM.
@@
Let Assumptions 2.1 (Differentiability), 2.2 (Convexity), 2.3 (Strong convexity of $f$), 2.8 (Star Similarity), and 6.1 hold. If the stepsize satisfies $\gamma\leq\frac{\mu}{4\delta_{*}^{2}}$, then SPPM satisfies, for every $k\geq 0$,
@@
The result shows linear convergence to a neighborhood of the solution, with the neighborhood size proportional to the variance $\sigma_{*}^{2}$ from Assumption 6.1, and depends on the stepsize $\gamma$. Using a decaying stepsize schedule leads to sublinear convergence to the exact solution.
We now present the convergence guarantee for the inexact variant of SPPM.
@@
Let Assumptions 2.1 (Differentiability), 2.2 (Convexity), 2.8 (Star Similarity), and 6.1 (Bounded Variance) hold. Consider SPPM-inexact with Assumption 2.9 satisfied. If the stepsize satisfies $\gamma\leq\frac{\mu(1-c)}{4\delta_{*}^{2}}$ and $T$ is chosen sufficiently large such that $\frac{\eta\gamma^{2}}{T^{\alpha}}\leq c$, where $0<c<1$ is a constant, then SPPM-inexact satisfies, for every $k\geq 0$,
$\mathbb{E}\big[\|x_{k}-x_{*}\|^{2}\big]\leq\left(1-\frac{1}{2}\min\left(\frac{\gamma\mu}{2},1-c\right)\right)^{k}\|x_{0}-x_{*}\|^{2}+\max\left(\frac{2}{\gamma\mu},\frac{1}{1-c}\right)\frac{4\gamma^{2}\sigma_{*}^{2}}{(1-c)}.$
The result guarantees convergence to a neighborhood of the solution, with a neighborhood size strictly larger than in the exact case due to the approximation error $c$. Specifically, the radius of this neighborhood increases by a factor of at least $\frac{1}{1-c}$. Additionally, the contraction factor becomes worse (closer to $1$), and the allowable stepsize is reduced by a factor of (1c)(1-c). Hence, the convergence speed is slower compared to the exact case, but convergence remains guaranteed as long as the approximation error is controlled.
@@ 7 Experiments
In this section, we present numerical experiments conducted for the optimization problem in the finite-sum form (2), where the functions $f_{i}$ take the specific form $f_{i}(x)=a_{i}\|x\|^{2s},$ with $s\in\mathbb{N}\setminus\{1\}$ and $a_{i}\in\mathbb{R}^{+}$ for all $i=1,\dots,n$. Each function $f_{i}$ is $(L_{0},L_{1})$-smooth, with $L_{0}=2s$ and $L_{1}=2s-1$, and is convex.
7.1 Main Experiments
We begin with two main experiments. The first experiment investigates the effect of varying the maximum number of iterations for the inner solver, without enforcing a stopping condition based on the gradient norm. The purpose of this analysis is to demonstrate that if the inner solver fails to achieve sufficient accuracy in solving the proximal step, the method may either diverge or exhibit slower convergence. We conduct two variations of this experiment.
In the first variation, we consider functions with different values of $s$, ranging from $2$ to $20$, and run the algorithm with varying maximum iteration limits for the inner solver: $T=5,20,50,100,500$. The number of functions is set to $n=100$, and the dimension is $d=5$. For the second variation, we modify the original problem to ensure strong convexity of $f$, enabling us to verify that the convergence rate becomes linear in this setting. The modified problem (2) is defined as follows: $f_{i}(x)=a_{i}\|x\|^{2s}+\lambda\langle e_{i},x\rangle^{2},$ where $a_{i}\in\mathbb{R}^{+}$ for all $i=1,\dots,n$, $\lambda\in\mathbb{R}^{+}$, and $e_{i}$ is a unit vector with its $i$-th coordinate equal to one. In this case, each function $f_{i}$ is $(L_{0},L_{1})$-smooth, with $L_{0}=2s+2\lambda$ and $L_{1}=2s-1$, and convex. Furthermore, $f$ is $\frac{\lambda}{n}$-strongly convex. We set $n=d=100$, $\lambda=2$, and $s=2$, and conduct experiments with inner solver iteration limits of $1,10,15,100,200$.
In the second experiment, we analyze the performance of SPPM-inexact and compare it with stochastic gradient descent (SGD) using constant stepsizes. Each function in this experiment has the form $f_{i}(x)=a_{i}|x|^{\frac{2-\alpha_{i}}{1-\alpha_{i}}}$, each of which is individually $\alpha_{i}$-smooth, as shown by Chen et al. (2023). It is straightforward to verify that each $f_{i}$ is also $1>\beta>\alpha_{i}>0$-symmetric smooth, implying all functions are $\alpha=\max_{i=1\dots N}\alpha_{i}$-symmetric smooth. We set $N=1000$, with coefficients $a_{i}$ uniformly distributed in the range $[10,1000)$.
We run both SPPM-inexact and SGD with three constant stepsizes: $\gamma\in\{10^{-3},5\times 10^{-7},10^{-8}\}$. Figure 2 illustrates that for smaller stepsizes ($\gamma=5\times 10^{-7},10^{-8}$), both methods exhibit slow convergence, stagnating at suboptimal solutions within the number of iterations considered. However, for the larger stepsize ($\gamma=10^{-3}$), SPPM-inexact converges consistently to the optimal solution, whereas SGD diverges.
This experiment highlights an advantage of SPPM-inexact: its capability to utilize relatively larger constant stepsizes effectively, resulting in faster and stable convergence. On the other hand, constant stepsize SGD may require smaller stepsizes or careful tuning to achieve convergence. Hence, while the considered optimization scenario is not inherently challenging, it clearly demonstrates the practical benefits of the flexibility in step-size selection provided by SPPM-inexact.
Figure 2: Comparison between SPPM-inexact and SGD with constant stepsizes $\gamma$.
7.2 Further Experiments
In addition to the main comparisons, we further examine the robustness of SPPM-inexact with respect to step-size selection and initialization. These experiments complement the previous results by providing deeper insight into the practical behavior of the method.
7.2.1 Impact of Different Stepsizes
@@
We next investigate the impact of different stepsizes on the convergence of the proposed method. Specifically, we aim to demonstrate that the method converges for any positive stepsize, provided an appropriately chosen tolerance for the inner solver is used, and that larger stepsizes lead to faster convergence rates. To validate this, we analyze three different values of $s$, namely $s=2,3,4$. For each case, five different stepsizes are tested: $\gamma=0.1,1,10,100,1000$. The number of functions is set to $n=1000$, the dimension to $d=100$, and an inexact solver is employed with a stopping criterion based on the squared norm of the gradient, with an accuracy threshold of $10^{-12}$: $\|\nabla\Psi_{k}(\hat{x}_{k})\|^{2}\leq 10^{-12}.$ As shown in Figure 3, the method converges for all chosen stepsizes. While larger stepsizes accelerate convergence in terms of the number of iterations, they also increase the computational cost per iteration. These observations confirm our theoretical findings.
7.2.2 Dependence on the Initial Point